OpenStudy (anonymous):

carol is baking cookies and needs 2 and 1/4 cups of flour, 2/3 cup of white sugar, 1/4 cup of brown suga, and 1/16 cup of baking soda. use this information to answer the following question. - if the amount of white sugar were increased by a factor of 2 and 1/2, in cups, how much sugar would there be? A. 2 and 1/2 B. 2 and 3/5 C. 1 and 2/3 D. 1 and 3/5

OpenStudy (anonymous):

will fan and medal !

OpenStudy (anonymous):

i think all u have to do is add the two fractions together. ( 2/3 + 2 1/2)

OpenStudy (johnweldon1993):

No see I was all confused...this wording sucks

OpenStudy (johnweldon1993):

It asks for TOTAL sugar...so I was adding white and brown together and was like...umm...no answer

OpenStudy (johnweldon1993):

But I guess all they want is White sugar sooo \[\large \text{Total Sugar} = \text{White Sugar} + \text{Brown Sugar}\] White sugar = 2/3 cup If the white sugar is increased by a factor of 2 1/2 we have multiplication: \[\large\text{White sugar} = \frac{2}{3} \times \frac{5}{2} = \frac{10}{6} = 1\frac{2}{3}\]

OpenStudy (johnweldon1993):

That top line was my original thinking...the total = white + brown...but just ignore that

OpenStudy (anonymous):

i just copied and pasted it from my work lol i know its confusing theres like 3 more problems and im pellet with fractions

OpenStudy (johnweldon1993):

Well look at what I did and let me know if anything doesn't make sense :)

OpenStudy (anonymous):

all u have to do is add the amount of white sugar it says then add what it tells u to... i dont get why ur multiplying...?

OpenStudy (johnweldon1993):

"increased by a factor of 2 1/2" increasing by a factor means multiplication :)

OpenStudy (anonymous):

i feel dumb lol

OpenStudy (johnweldon1993):

Dont...not at all lol fractions are confusing for many people so dont feel dumb :)

OpenStudy (anonymous):

can u answer another 1 or 2...?

OpenStudy (johnweldon1993):

...perhaps lol

OpenStudy (anonymous):

if the amount of flower were increased by a factor of 2 and 1/2 in cups how much flour would there be..? the original amount of flower is 2 and 1/4

OpenStudy (johnweldon1993):

Okay so same here original = 2 1/4 Increased by a factor of 2 1/2 So \[\large \frac{9}{4} \times \frac{5}{2} = ?\]

OpenStudy (anonymous):

sorry i had to takemy dog outside

OpenStudy (anonymous):

5 and 5/8 ..?

OpenStudy (johnweldon1993):

Perfect!

OpenStudy (anonymous):

yayy! okay, if only 1/4 the amount of brown sugar were used how much brown sugar would be used ..?

OpenStudy (johnweldon1993):

What is 1/4 of 1/4? \[\large \frac{1}{4} \times \frac{1}{4} = ?\]

OpenStudy (anonymous):

1/16..?

OpenStudy (johnweldon1993):

correct again :)

OpenStudy (anonymous):

last one lol! if only 1/4 the amount of baking soda in the recipe were used, how much baking soda was used..?

OpenStudy (johnweldon1993):

What is 1/4 of 2 1/4? \[\large \frac{1}{4} \times \frac{9}{4} = ?\]

OpenStudy (anonymous):

2 and 1/5..?

OpenStudy (anonymous):

wait thats wrong

OpenStudy (anonymous):

9/16 ii added the first time lol

OpenStudy (johnweldon1993):

Lol there ya go :)

OpenStudy (anonymous):

can that be simplified...?

OpenStudy (johnweldon1993):

Well are there any numbers that go into by 9 and 16 evenly?

OpenStudy (anonymous):

cause 9/16 isnt one of my choices lol

OpenStudy (anonymous):

my choices are A. 1/64 B. 1/20 C. 1/16 D. 1/4

OpenStudy (johnweldon1993):

Oh omg I used the flour not the baking soda XDD sorry about that \[\large \frac{1}{4} \times \frac{1}{16}\]

OpenStudy (anonymous):

its fine lol and sorry my laptop decided to restart its self

OpenStudy (anonymous):

1/64

OpenStudy (johnweldon1993):

there ya go...and no im sorry lol

OpenStudy (anonymous):

haha its fine i messed up to !